26.8 of Na 2 SO 4 × xH 2 O gave 12.6 g of water on heating. The value of x in the compound is (M. mass of Na 2 SO 4 = 142) (A) 14 g atoms of nitrogen (B) 0.6 moles of Sg (C) 4 g equivalents of sulphate ions (D) 1.2×10 24 chloride ions.



26.8 of Na2SO4 × xH2O gave 12.6 g of water on heating. The value of x in the compound is (M. mass of Na2SO4 = 142)


(A)        14 g atoms of nitrogen


(B)        0.6 moles of Sg


(C)        4 g equivalents of sulphate ions


(D)        1.2×1024 chloride ions.


Grade:11

1 Answers

Shane Macguire
10 years ago

Wt. of anhy. Na2SO4 = 26.8 – 12.6 = 14.2 g

 

        14.2 g anhy. Salt combines with water = 12.6 g

 

        142 g anhy. Salt combines with water

 

 

 

        \      H2O (moles) = 126/18 = 7